OpenStudy (anonymous):

7.04] Choose the correct simplification of 4 to the 3rd power over 4 to the 2nd power

jhonyy9 (jhonyy9):

4^3 ---- = 4^2 correct ?

jhonyy9 (jhonyy9):

so you need using x^a ---- = x^(a-b) x^b hope this will help you

OpenStudy (anonymous):

so it would be 1/4

jhonyy9 (jhonyy9):

no

OpenStudy (anonymous):

JUST 4

jhonyy9 (jhonyy9):

yes just 4 because there will get 4^(3-2)=4^1 =4 ok ?

OpenStudy (anonymous):

O OK I GET IT SO U JUST LEAVE X AS IT IS YOU DONT NEED TO CHANGE IT TO 1..THANK U

jhonyy9 (jhonyy9):

yw was my pleasure good luck bye